# Towards Backdoor Stealthiness in Model Parameter Space

### v1

@@ B.2. Different Architectures with Different Surrogate Models
B.3. Adversarial Backdoor Injection Does Not Impact Backdoor Effectiveness in Case of No Defense.
@@
B.4. Further TAC analysis
For clearer demonstration, we also provide sorted TAC value plots in Figure 9, which sorts the TAC values in Figure 2. Figure 9 demonstrates the existence of prominent neurons, and Grond is more stealthy.
@@ B.4. Further TAC analysis
B.5. Examples of Poisoned Images
Figure 10 shows four training images from ImageNet200 when applied with UPGD. Please note the images are only meant to demonstrate imperceptible trigger and poisoning perturbations. In our experiments, we only poison training images from the class “stingray” to inject the backdoor. The first row depicts poisoned images, while the second contains clean ones. Finally, the third row contains the residual of the first two rows. Notice that Grond does not introduce any visible difference to the clean images.

### v2

@@ C.4. Backdoor analysis by feature space inversion
C.6. Adversarial Backdoor Injection Does Not Impact Backdoor Effectiveness in Case of No Defense.
@@
C.7. Further TAC analysis
For a clearer demonstration, we also provide sorted TAC value plots in Figure 11, which sort the TAC values in Figure 2. Figure 11 demonstrates the existence of prominent neurons, and Grond is more stealthy.
@@ C.7. Further TAC analysis
Table 18. Evaluation of Grond against adversarial training (AT) on CIFAR10. RA refers to accuracy under PGD (Madry et al., 2018) adversarial attack.
Arch No Defense AT
BA ASR RA BA ASR RA
ResNet18 93.43 98.04 3.82 88.56 89.69 62.24
VGG16 92.57 90.10 9.64 85.64 85.78 59.75
DenseNet121 92.38 81.07 9.72 86.67 64.30 58.66
Figure 12. The weight changes after FT-SAM (Zhu et al., 2023) for baseline attacks.
C.8. Adversarial Training as A Defense
We evaluate Grond with adversarial training (AT), which is considered one of the most successful defenses against adversarial attacks, since the UPGD trigger is a special type of adversarial perturbation. Table 18 shows the backdoor performance and robustness of Grond before and after adversarial training using CIFAR10 with three architectures (ResNet18, VGG16, and DenseNet121). The robust accuracy against PGD attacks significantly increased after AT, confirming AT’s effectiveness. However, AT does not decrease ASR, which means Grond backdoor can resist AT.
C.9. Weight Changes After FT-SAM with 10 Baseline Attacks
Figure 12 provides the results of weight changes after applying FT-SAM (Zhu et al., 2023) to ten baseline attacks. The weight changes in baseline attacks show a significant difference in a few prominent neurons.
C.10. Attempts to Craft Inputs that Minimize Changes in the Parameter Space
In this section, we also try to craft poisoned training data that minimizes changes in the parameter space instead of using ABI.
In the previous design of Grond, the backdoor is injected by the UPGD trigger. Then, Grond, through ABI, limits the changes of backdoor neurons by reducing their weights. We consider replacing the ABI component with an operation only in the training data, which does not require control over training. Our goal is to partially break the effectiveness of injecting the backdoor via the poisoned data. Specifically, we add poisoned data into the target class with a poisoning rate of $PR=5\%$ under the clean-label setting, following the standard setting in previous work (Wang et al., 2023; Xu et al., 2024b). Then, we add poisoned data into every class with a poisoning rate of $PR\times\lambda$. We aim to use the poisoned data in non-targeted classes to partially break the effectiveness of poisoned data in the targeted class. Table 19 provides the experimental results with different $\lambda$ values. With the decrease in $\lambda$, both ASR (no defense) and ASR (after applying CLP) increase. This means that our attempt to minimize changes in the parameter space for stealthiness works. However, ASR after applying CLP is still lower than 10%, which means the attack is not robust against the backdoor defense.
Table 19. Crafting inputs to minimize changes in the parameter space.
$\lambda$ No Defense CLP
BA ASR BA ASR
0.5 94.09 1.43 93.39 1.43
0.1 94.23 6.09 93.43 1.97
0.01 93.69 26.83 90.90 3.87
0.001 93.88 68.81 92.80 4.89
0.0005 94.13 90.78 92.81 5.84
C.11. Examples of Poisoned Images
Figure 13 shows four training images from ImageNet200 when applied with UPGD. Please note that the images are only meant to demonstrate imperceptible trigger and poisoning perturbations. In our experiments, we only poison training images from the class “stingray” to inject the backdoor. The first row depicts poisoned images, while the second contains clean ones. Finally, the third row contains the residual of the first two rows. Notice that Grond does not introduce any visible difference to the clean images.